OpenStudy (anonymous):

write an equation of a line that passes through the points (2,3) and (2,-6)

OpenStudy (anonymous):

@Asexybabe17

OpenStudy (tkhunny):

You seem to be in the habit of posting questions without showing any work of your own. It would be good for you to change this behavior. SHOW YOUR WORK. Ever hear of the "2-Point Form" of a line?

rvc (rvc):

you did a similar q before

OpenStudy (anonymous):

I don't have any work!

OpenStudy (anonymous):

@Asexybabe17

OpenStudy (anonymous):

From two point form of the line: \[\frac{y - y1 }{ x - x1 } = \frac{ y2 - y1 }{x2 - x1 } \] \[\frac{ y - 3 }{ x - 2 } = \frac{ -6 - 3 }{ 2 - 2 }\] \[x = 2\] is the equation of the line. Notice that the abscissa of the points are the same i.e 2 So the line is a vertical line and its equation can be directly written by inspection. So the equation of the line is x = 2.

OpenStudy (tkhunny):

"I don't have any work! " Here's the deal... Get Some! You can't possibly have NO IDEA. Did you attend class AT ALL? Probably. Have you looked at your course materials AT ALL? Probably. Have you done any problem even remotely like this, EVER? Probably. Show us SOMETHING. Calculate a slope. Anything. I just don't believe you have NOTHING. I suspect strongly that you don't believe it, either. You teacher or instructor also doesn't believe it or you would not have been given this problem to work on. Show your work. No excuses.

OpenStudy (anonymous):

well you are wrong about 2 of those! because this is my 1st time having these kind of problems and there wasnt anything for me too look at about it!

OpenStudy (anonymous):

I have had this same paper for 2 months! I still haven't been able to figure them out!

OpenStudy (tkhunny):

I can't be wrong on any of those. They are listed a probabilities. 1) Stop arguing with me. 2) Stop making excuses. 3) Show your work. Do you know how to calculate a Slope? \(Slope = \dfrac{\Delta y}{\Delta x} = \dfrac{Change\;in\;y}{Change\;in\;x} = \dfrac{Rise}{Run} = \dfrac{y_{1} - y_{0}}{x_{1}-x_{0}} = m\) It has many names and descriptions. Have you ever seen any of these and can you now calculate such a thing with the given information?

OpenStudy (anonymous):

no i do not